# Fully stabilized Minkowski vacua in the $2^{6}$ Landau-Ginzburg model

### v1

@@ Abstract
We study moduli stabilization via fluxes in the $2^{6}$ Landau-Ginzburg model. Fluxes not only give masses to scalar fields but can also induce higher order couplings that stabilize massless fields. We investigate this for several different flux choices in the $2^{6}$ model and find two examples that are inconsistent with the Refined Tadpole Conjecture. We also present, to our knowledge, the first 4d $\mathcal{N}=1$ Minkowski solution in string theory without any flat direction.
@@ 1 Introduction
An intriguing way to avoid the complication of having to deal with a combination of fluxes and non-perturbative effects was proposed in Becker:2006ks . The mirror dual of a rigid CY-manifold has $h^{1,1}=0$ and therefore no Kähler moduli at all. That means one can stabilize all moduli in type IIB by using fluxes only. While there is no geometric interpretation of the internal space in this case, one can study such setups using the language of Landau-Ginzburg (LG) models Vafa:1988uu ; vafaOrbifoldized ; Witten:1993yc ; Candelas:1993nd . This was done almost twenty years ago in Becker:2006ks ; Becker:2007dn ; Becker:2007ee and has been recently revisited and extended in Ishiguro:2021csu ; Bardzell:2022jfh ; Becker:2022hse ; Cremonini:2023suw ; Becker:2023rqi ; Becker:2024ijy ; Ishiguro:2024coq . This allows one to explore flux compactifications away from the large volume, large complex structure, and even away from the weak coupling limit and thus provides a very interesting new place to test swampland conjectures.
@@
The Tadpole Conjecture states that the fluxes used to stabilize complex structure moduli contribute to the tadpole an amount $N_{\rm flux}$ that grows unacceptably fast with the number of moduli that are massive (or get stabilized). In particular, it is conjectured to be true that
$N_{\rm flux}>\frac{1}{3}n_{\rm stab}\,.$ (1)
For supersymmetric Minkowski vacua, this provides an upper bound on $n_{\rm stab}$, which is the number of moduli we can stabilize. This is due to the O3-plane tadpole cancellation condition
@@
where $N_{O3}$ is the finite number of O3-planes and $N_{D3}\geq 0$ the number of D3-branes.
The Tadpole Conjecture received a lot of attention recently and has been tested in many setups Bena:2021wyr ; Marchesano:2021gyv ; Plauschinn:2021hkp ; Lust:2021xds ; Bena:2021qty ; Grana:2022dfw ; Tsagkaris:2022apo ; Lust:2022mhk ; Coudarchet:2023mmm ; Braun:2023pzd ; Braun:2023edp ; Becker:2024ijy . These tests have been performed in the asymptotics as well as the interior of moduli space. However, none of these papers found a violation of the Tadpole Conjecture. The $2^{6}$ LG model that we study in this paper, however, allows for violations of the conjecture as we will explain in this paper. This result was independently obtained in Becker2024 , where the authors initiate a full classification of all flux configurations in this model.
While the Tadpole Conjecture is mostly relevant for compactifications with a large number of complex structure moduli, it was independently conjectured that all 10d supergravity compactifications to 4d Minkowski solutions should always have a massless scalar field Andriot:2022yyj . This so-called Massless Minkowski Conjecture is satisfied in all known supergravity compactifications but violated in this non-geometric $2^{6}$ LG model Becker2024 . The existence of 4d $\mathcal{N}=1$ vacua without massless fields is extremely interesting and shows how useful these non-geometric LG models are in improving our understanding of the string theory landscape of vacua.
@@ 2.1 Homology and Cohomology
where we have defined the conjugate index notation $\bar{l}:=4-l$. The astute reader may have noticed that the factor of $4$ is switched in equations (20) and (21) compared to the conventions in the recent paper Becker:2024ijy on the $1^{9}$ LG model. The present choice is made following the conventions of [^1] The concurrently published paper Becker2024 also adopts the same normalization as ours. Becker:2006ks . We will say more on this soon.
@@ 3.2 One to three $\Omega$ flux choices
Before ending this subsection we state the super important result that the flux choice $G^{4}$ has $N_{\rm flux}=16$ and 52 massive fields. To our knowledge, this is the first instance in the literature that violates the tadpole bound: $N_{\rm flux}>\tfrac{1}{3}n_{\rm stab}$. Thus the $2^{6}$ LG model seems to violate the Refined Tadpole Conjecture in its original form Bena:2020xrh . [^8] The tadpole condition in this example is not satisfied and requires the presence of 24 D3-branes. Those might have moduli associated with themselves but that is irrelevant for the fact that 52 complex structure moduli are massive for this simple flux choice. We comment on this further in section 4 below.
@@ 3.3 A four $\Omega$ flux choice
Given the slightly disappointing outcome for the stabilization at higher order for the simplest solutions in the $2^{6}$ model, we now turn to an intermediate solution that is much more interesting. This flux choice was discovered by the authors of Becker2024 and they kindly allowed us to study the higher-order stabilization for it, already before they published their paper. The flux choice is
@@
Solving the above constraints leads to multiple branches of solutions. There are branches which fix $8$, $9$, $10$, and $12$ fields. The branches that stabilize $9,10$ and, $12$ fields already violate the bound of the Refined Tadpole Conjecture. However, the branches that stabilize $8$ fields do not violate the bound. As explained in Becker:2024ijy , when such branches with different numbers of stabilized fields appear, we pick the branch with the least number of stabilized fields. So at cubic order, we do not have a violation of the bound yet. We then analyse the non-trivial cubic constraints arising after the inclusion of quartic terms in the superpotential. The minimum number of stabilized fields occur on branches that had $8$ or $9$ fields stabilized at cubic order. There are branches that had stabilized $9$ fields at cubic order, that stabilize no fields at quartic or quintic order, leaving us with $75+9+0+0=84=28\cdot 3$ stabilized fields. So, this would violate the original Refined Tadpole Conjecture with $\alpha<1/3$. All branches that stabilized $8$ fields at cubic order stabilize at least $1$ field at quartic order, thus leading to $75+8+1=84=28\cdot 3$ and the same conclusion as above.
@@ 3.4 Close to Full Rank Solutions
For the curious reader, let us briefly mention how we obtained these high rank solutions. There is a straightforward algorithm to obtain an integral basis of the supersymmetric flux lattice in the present model. It is the same algorithm that was used in Becker:2023rqi to derive an integral basis for Minkowski fluxes in the $1^{9}$ model. See also Becker:2024ijy for a detailed description. We employed this algorithm in the $2^{6}$ model to obtain $180$ basis elements for the Minkowski lattice. Using “small linear combinations” – with four or five basis vectors linearly combined with small integer coefficients (between $-2$ and $2$) – it was possible to generate solutions with a large variety of $(N_{\rm flux},n_{\rm massive})$ values. All of our high rank solutions were obtained this way, with a remarkably small computational effort. The authors of Becker2024 , in their attempt to extensively test the tadpole conjecture, use this as one of the methods to generate a large number of solutions.
@@ 3.5 Fully stabilized Minkowski vacua
Lastly, we present two solutions that were part of our randomly generated set of solutions with $N_{\rm flux}=40$ and that actually do stabilize all fields through higher order terms. Given that there are only 2 out of 142 examples, this shows that these are rare but they do exist. These two solutions provide to our knowledge the first Minkowski vacua without any flat directions (besides 11d supergravity). They both have 86 massive fields, 0 stabilized fields when including cubic terms via $W_{3}$, and finally 5 stabilized fields when including $W_{4}$. At that point, we do not need to worry about higher-order terms since we cannot stabilize any more fields and $\partial_{i}W_{n}=0$ is automatically satisfied for all $n$ since all moduli are stabilized at the origin of moduli space. Note that these solutions have $N_{\rm flux}=40$ and 91 stabilized fields. They are thus not in tension with the Refined Tadpole Conjecture that would allow the stabilization of up to 119 fields for $N_{\rm flux}=40$.
@@
In a paper that appeared at the same time as ours Becker2024 , the authors found many flux configurations in this model that have $N_{\rm flux}=40$ and full rank 91 for the mass matrix. So, these solutions in Becker2024 are the first Minkowski vacua in which all scalar fields are massive. We comment further on the significance of these and our solutions in the next section.
@@ 4.1 Full moduli stabilization and 4d $\mathcal{N}=1$ vacua
Finding new simple models with full moduli stabilization has thus become an important goal in string phenomenology. Non-geometric LG models provide an intriguing way of doing so. On the one hand, they allow us to bring the full strength of string theory to the problem. On the other hand, they give rise to very simple 4d $\mathcal{N}=1$ supergravity theories that depend, in the presence of fluxes, on all moduli. The initial study of the simplest of these models, the $1^{9}$ model, has discovered the presence of many massless fields and potential directions that do not get lifted even at higher order Becker:2024ijy . However, the next more complicated example, the $2^{6}$ model that is studied here and at the same time in Becker2024 has completely changed the picture. Fully stabilized, isolated Minkowski vacua have been found in this paper albeit with some massless fields. The authors of Becker2024 found isolated Minkowski vacua without massless fields. Thus, there are now new constructions of moduli stabilization for the community to understand better and scrutinize in the future.
Let us already provide here a few comments on why we believe this moduli stabilization scenario is trustworthy. We have solved the equations $W=\partial_{i}W=0$, which do not depend on the Kähler potential at all but only on the superpotential $W$. For non-geometric LG models it was shown in Becker:2006ks that the superpotential is exact and does not receive any perturbative or non-perturbative corrections. For example, the absence of the standard ED3 instanton corrections follows from $h^{1,1}=0$, i.e., the absence of 4-cycles. For D(-1) instantons it was likewise argued that they vanish in these models. This is consistent with the recent results in Kim:2022jvv , where it was shown that they are only non-zero in models where the D7-brane charge is not canceled locally. Since the LG models have no D7-branes or O7-planes there is no such charge at all. Thus, the superpotential does not receive any corrections at all and any Minkowski vacuum is actually trustworthy. This reasoning applies to the two solutions above in equations (91) and (92) that have 86 massive and 5 massless but stabilized scalar fields. It also applies equally to the Minkowski vacua found in the paper Becker2024 that have 91 massive fields and therefore do not require any stabilization at higher order.
@@ 4.2 The Massless Minkowski Conjecture
The Massless Minkowski Conjecture states that 10d supergravity compactifications to 4d Minkowski space always have a massless scalar field. Here we are not dealing with compactifications of type IIB supergravity so we might be violating an important assumption of the conjecture. So, strictly speaking, the LG models cannot violate this conjecture. However, let us nevertheless entertain the possibility that all Minkowski vacua have a massless scalar field.
@@
As we discussed in the previous subsection, there are no perturbative or non-perturbative corrections for these non-geometric LG models. Thus, due to these powerful non-renormalization theorems, we do not have to worry about infinite numbers of corrections and can trust the fully stabilized Minkowski vacua we found in this paper. This reasoning of course also applies to the Minkowski vacua found in the paper Becker2024 that have 91 massive fields and do not even require any stabilization at higher order.
@@ 4.3 The Tadpole Conjecture
$N_{\rm flux}=\frac{1}{\tau-\bar{\tau}}\int G_{3}\wedge\bar{G}_{3}>\alpha\,n_{\rm stab}\,,$ (101)
@@
$N_{\rm flux}=\frac{1}{\tau-\bar{\tau}}\int G_{3}\wedge\bar{G}_{3}>\alpha\,n_{\rm massive}\,,$ (102)
@@
Either form of the conjecture is violated by our flux choice above in equation (57). That flux choice had $N_{\rm flux}=16$ and 52 massive fields, which gives
@@
However, the authors of Becker:2024ijy generated a large set of flux configurations in the $1^{9}$ LG model and found a spectacular confirmation of the Tadpole Conjecture (see figure 2 on page 22 of that paper). A similar study for the $2^{6}$ LG model was undertaken in the paper Becker2024 that appeared at the same time as ours. The authors kindly shared their results with us and those do clearly show a confirmation of a bound like the one in equation (101) albeit with a smaller coefficient. The authors of Becker2024 propose $\alpha=1/4$ and it could be that this is the correct statement.

### v2

@@ Abstract
We study moduli stabilization via fluxes in the $2^{6}$ Landau-Ginzburg model. Fluxes not only give masses to scalar fields but can also induce higher order couplings that stabilize massless fields. We investigate this for several different flux choices in the $2^{6}$ model and find two examples that are violating the Refined Tadpole Conjecture by more than a factor of 2. We also present, to our knowledge, the first 4d $\mathcal{N}=1$ Minkowski solution in string theory without any flat direction.
@@ 1 Introduction
An intriguing way to avoid the complication of having to deal with a combination of fluxes and non-perturbative effects was proposed in Becker:2006ks . The mirror dual of a rigid CY-manifold has $h^{1,1}=0$ and therefore no Kähler moduli at all. That means one can stabilize all moduli in type IIB by using fluxes only. While there is no geometric interpretation of the internal space in this case, one can study such setups using the language of Landau-Ginzburg (LG) models Vafa:1988uu ; vafaOrbifoldized ; Schimmrigk:1992ai ; Witten:1993yc ; Candelas:1993nd . This was done almost twenty years ago in Becker:2006ks ; Becker:2007dn ; Becker:2007ee and has been recently revisited and extended in Ishiguro:2021csu ; Bardzell:2022jfh ; Becker:2022hse ; Cremonini:2023suw ; Becker:2023rqi ; Becker:2024ijy ; Ishiguro:2024coq . This allows one to explore flux compactifications away from the large volume, large complex structure, and even away from the weak coupling limit and thus provides a very interesting new place to test swampland conjectures.
@@
The Tadpole Conjecture states that the fluxes used to stabilize complex structure moduli contribute to the tadpole an amount $N_{\rm flux}$ that grows unacceptably fast with the number of moduli that are massive (or get stabilized). In particular, it is conjectured to be true that [^1] The first version of this paper and previous papers studying moduli stabilization in the $1^{9}$ LG model Becker:2022hse ; Becker:2024ijy , did not have the factor of 2 in the equation below. This factor of 2 arises from us defining the flux tadpole in equation (34) in the covering space following Becker:2006ks . The original Tadpole Conjecture paper Bena:2020xrh has in equation (2.1) a factor of 1/2 in front of the flux contribution, effectively counting the flux contribution in the quotient space. We thank Daniel Junghans for bringing this to our attention.
$N_{\rm flux}>2\,\alpha\,n_{\rm stab}\,.$ (1)
where the Refined Tadpole Conjecture states that $\alpha=1/3$. For supersymmetric Minkowski vacua, this provides an upper bound on $n_{\rm stab}$, which is the number of moduli we can stabilize. This is due to the O3-plane tadpole cancellation condition
@@
where $N_{O3}$ is the finite number of O3-planes and $N_{D3}\geq 0$ the number of D3-branes, both counted in the covering space.
The Tadpole Conjecture received a lot of attention recently and has been tested in many setups Bena:2021wyr ; Marchesano:2021gyv ; Plauschinn:2021hkp ; Lust:2021xds ; Bena:2021qty ; Grana:2022dfw ; Tsagkaris:2022apo ; Becker:2022hse ; Lust:2022mhk ; Coudarchet:2023mmm ; Braun:2023pzd ; Braun:2023edp ; Becker:2024ijy . These tests have been performed in the asymptotics as well as the interior of moduli space. However, none of these papers found a violation of the Tadpole Conjecture except Becker:2022hse ; Becker:2024ijy . These two papers that studied the $1^{9}$ LG model, confirmed the linear relationship proposed by the Tadpole Conjecture, but found a violation of the Refined Tadpole Conjecture by almost a factor of 2. The $2^{6}$ LG model that we study in this paper allows for violations of the Refined Tadpole Conjecture by more than a factor of 2, as we will explain in this paper. This result was independently obtained in Becker:2024ayh , where the authors initiate a full classification of all flux configurations in this model.
While the Tadpole Conjecture is mostly relevant for compactifications with a large number of complex structure moduli, it was independently conjectured that all 10d supergravity compactifications to 4d Minkowski solutions should always have a massless scalar field Andriot:2022yyj . This so-called Massless Minkowski Conjecture is satisfied in all known supergravity compactifications but violated in this non-geometric $2^{6}$ LG model Becker:2024ayh . The existence of 4d $\mathcal{N}=1$ vacua without massless fields is extremely interesting and shows how useful these non-geometric LG models are in improving our understanding of the string theory landscape of vacua.
@@ 2.1 Homology and Cohomology
where we have defined the conjugate index notation $\bar{l}:=4-l$. The astute reader may have noticed that the factor of $4$ is switched in equations (20) and (21) compared to the conventions in the recent paper Becker:2024ijy on the $1^{9}$ LG model. The present choice is made following the conventions of [^2] The concurrently published paper Becker:2024ayh also adopts the same normalization as ours. Becker:2006ks . We will say more on this soon.
@@ 3.2 One to three $\Omega$ flux choices
Before ending this subsection we state the important result that the flux choice $G^{4}$ has $N_{\rm flux}=16$ and 52 massive fields. To our knowledge, this is the first instance that violates the Refined Tadpole Conjecture bound Bena:2020xrh by more than a factor of 2: $N_{\rm flux}>\tfrac{2}{3}n_{\rm stab}$. [^9] The tadpole condition in this example is not satisfied and requires the presence of 24 D3-branes. Those might have moduli associated with themselves but that is irrelevant for the fact that 52 complex structure moduli are massive for this simple flux choice. We comment on this further in section 4 below.
@@ 3.3 A four $\Omega$ flux choice
Given the slightly disappointing outcome for the stabilization at higher order for the simplest solutions in the $2^{6}$ model, we now turn to an intermediate solution that is much more interesting. This flux choice was discovered by the authors of Becker:2024ayh and they kindly allowed us to study the higher-order stabilization for it, already before they published their paper. The flux choice is
@@
Solving the above constraints leads to multiple branches of solutions. There are branches which fix $8$, $9$, $10$, and $12$ fields. The branches that stabilize $9,10$ and, $12$ fields already violate the bound of the Refined Tadpole Conjecture by more than a factor of 2, which did not seem possible in the $1^{9}$ LG model Becker:2024ijy . However, the branches that stabilize $8$ fields do not violate the bound by that much. As explained in Becker:2024ijy , when such branches with different numbers of stabilized fields appear, we pick the branch with the least number of stabilized fields. So at cubic order, we do not have a violation of the bound by a factor of 2 yet. We then analyse the non-trivial cubic constraints arising after the inclusion of quartic terms in the superpotential. The minimum number of stabilized fields occur on branches that had $8$ or $9$ fields stabilized at cubic order. There are branches that had stabilized $9$ fields at cubic order, that stabilize no fields at quartic or quintic order, leaving us with $75+9+0+0=84=28\cdot 3$ stabilized fields. So, this would violate the original Refined Tadpole Conjecture by a factor of 2. All branches that stabilized $8$ fields at cubic order stabilize at least $1$ field at quartic order, thus leading to $75+8+1=84=28\cdot 3$ and the same conclusion as above.
@@ 3.4 Close to Full Rank Solutions
For the curious reader, let us briefly mention how we obtained these high rank solutions. There is a straightforward algorithm to obtain an integral basis of the supersymmetric flux lattice in the present model. It is the same algorithm that was used in Becker:2023rqi to derive an integral basis for Minkowski fluxes in the $1^{9}$ model. See also Becker:2024ijy for a detailed description. We employed this algorithm in the $2^{6}$ model to obtain $180$ basis elements for the Minkowski lattice. Using “small linear combinations” – with four or five basis vectors linearly combined with small integer coefficients (between $-2$ and $2$) – it was possible to generate solutions with a large variety of $(N_{\rm flux},n_{\rm massive})$ values. All of our high rank solutions were obtained this way, with a remarkably small computational effort. The authors of Becker:2024ayh , in their attempt to extensively test the tadpole conjecture, use this as one of the methods to generate a large number of solutions.
@@ 3.5 Fully stabilized Minkowski vacua
Lastly, we present two solutions that were part of our randomly generated set of solutions with $N_{\rm flux}=40$ and that actually do stabilize all fields through higher order terms. Given that there are only 2 out of 142 examples, this shows that these are rare but they do exist. These two solutions provide to our knowledge the first Minkowski vacua without any flat directions (besides 11d supergravity). They both have 86 massive fields, 0 stabilized fields when including cubic terms via $W_{3}$, and finally 5 stabilized fields when including $W_{4}$. At that point, we do not need to worry about higher-order terms since we cannot stabilize any more fields and $\partial_{i}W_{n}=0$ is automatically satisfied for all $n$ since all moduli are stabilized at the origin of moduli space. Note that these solutions have $N_{\rm flux}=40$ and 91 stabilized fields. They are thus violating the Refined Tadpole Conjecture that would allow only the stabilization of less than $6040\cdot\frac{3}{2}=$ 60 fields for $N_{\rm flux}=40$.
@@
In a paper that appeared at the same time as ours Becker:2024ayh , the authors found many flux configurations in this model that have $N_{\rm flux}=40$ and full rank 91 for the mass matrix. So, these solutions in Becker:2024ayh are the first Minkowski vacua in which all scalar fields are massive. We comment further on the significance of these and our solutions in the next section.
@@ 4.1 Full moduli stabilization and 4d $\mathcal{N}=1$ vacua
Finding new simple models with full moduli stabilization has thus become an important goal in string phenomenology. Non-geometric LG models provide an intriguing way of doing so. On the one hand, they allow us to bring the full strength of string theory to the problem. On the other hand, they give rise to very simple 4d $\mathcal{N}=1$ supergravity theories that depend, in the presence of fluxes, on all moduli. The initial study of the simplest of these models, the $1^{9}$ model, has discovered the presence of many massless fields and potential directions that do not get lifted even at higher order Becker:2024ijy . However, the next more complicated example, the $2^{6}$ model that is studied here and at the same time in Becker:2024ayh has completely changed the picture. Fully stabilized, isolated Minkowski vacua have been found in this paper albeit with some massless fields. The authors of Becker:2024ayh found isolated Minkowski vacua without massless fields. Thus, there are now new constructions of moduli stabilization for the community to understand better and scrutinize in the future.
Let us already provide here a few comments on why we believe this moduli stabilization scenario is trustworthy. We have solved the equations $W=\partial_{i}W=0$, which do not depend on the Kähler potential at all but only on the superpotential $W$. For non-geometric LG models it was shown in Becker:2006ks that the superpotential is exact and does not receive any perturbative or non-perturbative corrections. For example, the absence of the standard ED3 instanton corrections follows from $h^{1,1}=0$, i.e., the absence of 4-cycles. For D(-1) instantons it was likewise argued that they vanish in these models. This is consistent with the recent results in Kim:2022jvv , where it was shown that they are only non-zero in models where the D7-brane charge is not canceled locally. Since the LG models have no D7-branes or O7-planes there is no such charge at all. Thus, the superpotential does not receive any corrections at all and any Minkowski vacuum is actually trustworthy. This reasoning applies to the two solutions above in equations (91) and (92) that have 86 massive and 5 massless but stabilized scalar fields. It also applies equally to the Minkowski vacua found in the paper Becker:2024ayh that have 91 massive fields and therefore do not require any stabilization at higher order.
@@ 4.2 The Massless Minkowski Conjecture
The Massless Minkowski Conjecture states that 10d supergravity compactifications to 4d Minkowski space always have a massless scalar field Andriot:2022yyj . Here we are not dealing with compactifications of type IIB supergravity so we might be violating an important assumption of the conjecture. So, strictly speaking, the LG models cannot violate this conjecture. However, let us nevertheless entertain the possibility that all Minkowski vacua have a massless scalar field.
@@
As we discussed in the previous subsection, there are no perturbative or non-perturbative corrections for these non-geometric LG models. Thus, due to these powerful non-renormalization theorems, we do not have to worry about infinite numbers of corrections and can trust the fully stabilized Minkowski vacua we found in this paper. This reasoning of course also applies to the Minkowski vacua found in the paper Becker:2024ayh that have 91 massive fields and do not even require any stabilization at higher order.
@@ 4.3 The Tadpole Conjecture
$N_{\rm flux}=\frac{1}{\tau-\bar{\tau}}\int G_{3}\wedge\bar{G}_{3}>2\,\alpha\,n_{\rm stab}\,,$ (101)
@@
$N_{\rm flux}=\frac{1}{\tau-\bar{\tau}}\int G_{3}\wedge\bar{G}_{3}>2\,\alpha\,n_{\rm massive}\,,$ (102)
@@
Either form of the conjecture is violated by our flux choice above in equation (57) by more than a factor of 2, which seemed to be a previously found weaker bound in the $1^{9}$ LG model Becker:2024ijy . That flux choice in equation (57) had $N_{\rm flux}=16$ and 52 massive fields, which gives
@@
However, the authors of Becker:2024ijy generated a large set of flux configurations in the $1^{9}$ LG model and found a confirmation of the Tadpole Conjecture albeit with a different factor (see figure 2 on page 22 of that paper). A similar study for the $2^{6}$ LG model was undertaken in the paper Becker:2024ayh that appeared at the same time as ours. The authors kindly shared their results with us and those do show a confirmation of a bound like the one in equation (101) albeit with a much smaller coefficient. Interestingly, the authors of Becker:2024ayh find $\alpha\lesssim 1/6.8$ which is smaller than the F-theory expectation $\alpha=1/4$.